OpenStudy (anonymous):

divisor=3 cand=3 count=2 while(count<1000): if cand%divisor==0: if cand==divisor: cand=cand+2 divisor=3 count=count+1 print 'inside_if_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) else: cand=cand+2 divisor=3 print 'inside_else_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) else: divisor=divisor+1 print 'outside_else_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) prin

OpenStudy (anonymous):

I feel like i'm close to getting PS1, but once my counter hits 999, it hits the first inside_if_statement again and thus adds 2 to the 1000th prime. I'm not sure of a way around this... Would someone mind taking a look?

OpenStudy (microbot):

what is the last print statement? i think there is some of your code missing.

OpenStudy (anonymous):

my final print statement basically says "print cand"

OpenStudy (anonymous):

figured it out--here's the working version: divisor=3 cand=3 count=2 while(count<1001): if cand%divisor==0: if cand==divisor: cand=cand+2 divisor=3 count=count+1 ##print 'inside_if_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) if count==1001: cand=cand-2 else: cand=cand-1+1 else: cand=cand+2 divisor=3 ##print 'inside_else_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) else: divisor=divisor+1 ##print 'outside_else_statement counter '+str(count)+'cand '+str(cand)+'divisor '+str(divisor) print cand

OpenStudy (microbot):

im glad i could help :P xD now seriously : gj for figuring it out!:)and im sry i couldnt help cause i had no time:P maybe next time !